How much work is required to raise a 0.1 kg block to a height of 2 m and simultaneously give it a constant velocity of 3 m/s? b) What is the change in its potential energy? c) What is the power output of the upward force in watts?


1
Expert's answer
2021-05-27T18:38:02-0400

(a)"W=mgh+\frac{1}{2}mv^2"

"=(0.1\times9.8\times2)+(\frac{1}{2}\times0.1\times3^2)"

"=2.41J"

(b) "\triangle P=P_2-P_1 \space" but "P_1=0"

"\triangle P=mgh=0.1\times9.8 \times2=1.96J"

(c) "P=F\times V=0.1\times 9.8\times3=2.94W"